an oxygen cylinder of volume 2.3 cubic foot has a pressure of 2200 psi gauge and is 21 degrees Celsius. Determine the mass of oxygen in the cylinder



1
Expert's answer
2022-03-08T13:51:02-0500

The equation of state of a perfect gas says

"pV=\frac{m}{M}RT"

Hence, the mass of an oxygen

"m=pVM\/RT"

"m=\frac{(15.17*10^6\rm \: Pa)*(0.065\: m^3)*0.016 \: kg/mol}{(8.31\:\rm J/mol\cdot K)*294\:K}=6.46\:\rm kg"